OpenStudy (anonymous):

If f(x)= x^2 + 2x, what is f(x-1)? Explain please. a) x^2 - 3 b) x^2 - 1 c)x^2 + 1 d)x ^2 + 3

OpenStudy (dayakar):

just u have to do is replace x with x-1 and simplify

OpenStudy (whpalmer4):

it might be simpler to understand this if you rewrite it slightly. \[f(u)=u^2+2u\]is completely equivalent to your original. it is a recipe for calculating the value of the function \(f\). you take the input value (which may be a number or an algebraic expression). multiply it by itself, then add twice the input value. if we want to find \(f(x-1)\) that is the same as letting \(u=x-1\) in my version: \[f(x-1) = f(u) = u^2+2u = (x-1)^2+2(x-1)\] Now just expand and simplify the expression \((x-1)^2+2(x-1)\) and you have your answer. \[(x-1)^2+2x=(x-1)(x-1)+2x=x^2-x-x+1+2x-2=x^2-1\] it might be illustrative to try plugging in an actual number or two. \[f(3)= u^2+2u = 3^2+2*3=9+6=15\] but what if that \(3\) was written as \(3=4-1\), so we can use the \((x-1)\) form? \[f(3)=f(4-1) = (x-1)^2+2(x-1)=x^2-1=4^2-1=15\] same answer!